OpenStudy (mwhit2101):

SIXTH GRADE MATH HELP MEDAL AND FAN In the figures, Triangle CDA is congruent to Triangle XYZ. Name the congruent angles.

OpenStudy (mwhit2101):

Choice #1 /C=/D,/A, = /X, /Y= /Z Choice #2 /C=/X, /D= /Y, /A= /Z Choice #3 /C=/Z,/D= /Y, /A= /X

OpenStudy (anonymous):

#2

OpenStudy (mwhit2101):

How do you know

OpenStudy (anonymous):

read the problem..its like reading from left to right...C=1 D=2 A=3 then you have the second one..x=1 y=2 z=3

OpenStudy (mwhit2101):

oK thanks :)

OpenStudy (anonymous):

np